# Achieving Distributed MIMO Performance with Repeater-Assisted Cellular Massive MIMO

### v2

@@ Abstract
In what ways could cellular massive MIMO be improved? This technology has already been shown to bring huge performance gains. However, coverage holes and difficulties to transmit multiple streams to multi-antenna users because of insufficient channel rank remain issues. Distributed MIMO, also known as cell-free massive MIMO, might be the ultimate solution. However, while being a powerful technology, it is expensive to install backhaul, and it is a difficult problem to achieve accurate phase alignment for coherent multiuser beamforming on downlink. Another option is reflective intelligent surfaces – but they have large form factors and require a lot of training and control overhead, and probably, in practice, some form of active filtering to make them sufficiently band-selective.
We propose a new approach to densification of cellular systems, envisioning repeater-assisted cellular massive MIMO, where a large numbers of physically small and cheap wireless repeaters are deployed. They receive and retransmit signals instantaneously, appearing as ordinary scatterers in the channel but with amplification. We elaborate on the requirements of such repeaters, show that the performance of these systems could potentially approach that of distributed MIMO, and outline future research directions.
@@ I Introduction
Massive multiple-input multiple-output (MIMO) has become the 5G technology and will continue to be one of the cornerstones in cellular systems. Sub-6 GHz time division-duplex (TDD) multi-user massive MIMO systems with reciprocity-based beamforming can serve many terminals in the same time-frequency resource, also in high mobility scenarios. By phase-coherent combining, an array gain is achieved, which can significantly improve signal quality and link budget.
@@
Although the opportunities with D-MIMO systems are many, they impose deployment issues. Complex questions, such as how to get the data to the distributed units, where to perform the processing, how to synchronize the APs, and how to distribute the clocks and phase references remain unsolved. Furthermore, the main concepts in the standards are designed for co-located MIMO (C-MIMO) and going to D-MIMO would be a big step for industry to adopt. There would be a need for radical network architectural changes and a large increase in site and maintenance costs. In industry there can be some reluctance to big changes and smaller steps, each beneficial, would be more likely to be implemented.
To make the transition smoother and challenge the existing views on how densification of current cellular systems could be done, we envision repeater-assisted massive MIMO (RA-MIMO). A similar vision was presented in [1], but extra delay, increased delay spread and compromised channel reciprocity were identified as obstacles; these issues are solved in our vision. Repeaters have traditionally been deployed to improve coverage. Coverage improvements and capacity scaling in MIMO systems aided by repeaters was studied in LoS environments [2] and full-duplex radio relays were investigated to improve outdoor-to-indoor communication [3]. Today modern massive MIMO base stations enable multi-user MIMO operation. Challenging the current views on how densification could be done; we propose to rethink repeaters, could they be deployed inside the cell to support massive C-MIMO systems to increase macro diversity, acting as channel scatterers but with amplification, and potentially approach D-MIMO performance?
Note the distinction between a repeater and a relay in terms of the delay introduced. A repeater instantaneously amplifies and re-transmits the signal, with a delay of at most a few hundred nanoseconds. A relay receives a signal of substantially longer duration (e.g., half a slot) and then either amplifies and forwards, or decodes and forwards that signal. Repeaters act as additional channel scatterers with amplification, but relays do not.
@@
Other ways to enhance performance, without creating extra cells, are to add extra APs or, use reflective intelligent surfaces (RISs). Table I presents a comparison between these three options. A RIS could perform the work of a repeater. Network-controlled repeaters (NCRs) are under consideration in standardization activities [4], while RISs have been a topic of pre-standardization discussions. Benefits of deploying repeaters or RISs, compared to extra APs, are that no phase synchronization nor backhaul is needed.
Comparing repeaters and RISs, repeaters are easier to deploy due to their small form factor. Training or significant control signaling is not required. While bringing similar benefits, a RIS has a much larger form factor, requires significant control signaling, and incurs significant overhead for training and configuration. Furthermore, a passive, or semi-passive RIS would not be nearly as band-selective as a repeater, which can be frequency-selective down to the level of a single channel. An active RIS, which could amplify signals, could act similar as a repeater but due to regulatory requirements it would likely require a large number of expensive components to filter the signal. As a first-order approximation, in order to achieve the same signal amplification as a repeater, a [passive] RIS would require a number of unit cells equal to the square root of the repeater (power) gain. To replace a 60-dB gain repeater, a RIS would need in the order of 31 × 31 ~ 1000 unit cells (for simplicity, we ignore the directivity of the patch antenna patterns here). Furthermore, to obtain the macro diversity, one would need many distributed units, which appears challenging with RISs.
@@
Array gain None High Some
@@ II Envisioning repeater-assisted massive MIMO
The signal received at the base station or the terminal will be a superposition of signals traveling via direct paths, including LoS and/or multipath reflections, and signals going via repeater(s). This means that the channels between a terminal and the base station is a result of several components: 1) the direct path between the terminal and base station, 2) paths between a terminal and a repeater, 3) paths between a repeater and a scatterer, 4) paths between repeaters, 5) paths between a repeater and the base station, and 6) paths between a scatterer and the base station. As this is only concatenations of linear time-invariant (LTI) systems, the result is also an LTI system. Its response is (in principle) the same in uplink and downlink, enabling reciprocity-based multi-user beamforming. Repeaters should act as ordinary channel scatterer, but with higher gain. The repeaters will naturally introduce a delay, i.e. a frequency-dependent phase shift, which is typically not a problem as it is comparable to the frequency response of a far-away scatterer.
III Repeaters enabling the vision
@@ III-A Requirements on the repeater
To support the vision of RA-MIMO systems, the key aspect is that it must be possible to realize repeaters that have the same properties as ordinary channel scatterers present in the propagation environment. This makes them a transparent part from both the base station and terminal perspective. To act as a channel scatterer, the repeater must satisfy the following requirements:
@@
    Noise is unavoidable and stems from multiple sources: thermal noise, quantization noise, and phase noise. There is a limit on how noisy a repeater can be in order to not act as a limiting noise source.
@@ III-C Dual-antenna repeaters
Radio frequency (RF) band filters can be used, see 3) in Fig. 2, to attenuate out-of-band signals before the LNA, and to comply with out-of-band emission requirements. However, such filters may not be sufficient to fulfill gain requirements outside the repeater passband. Better selectivity can be achieved by baseband filtering, as shown in 4) in Fig. 2: the received signal is then down-converted to baseband, filtered, and up-converted using the same local oscillator (LO).
@@ IV Performance of repeater-assisted cellular massive MIMO
To show the potential of RA-MIMO, we compare it to massive C-MIMO, and densely deployed D-MIMO, which here serves as a benchmark for ideal performance, e.g. assuming no synchronization issues or other impairments. We consider multi-user MIMO transmission, where either a massive MIMO array located at the base station (first two cases) or distributed APs (last case) receive uplink data from multiple terminals, whereas the repeaters in RA-MIMO are located at the same positions as the D-MIMO APs. For simplicity, the results herein disregard interactions between repeaters, however, the maximum amplification gain is chosen such that the system becomes stable without positive feedback (using the criteria in [6]). The signal-to-interference-plus-noise ratio (SINR) after the uplink combiner is used as performance metric, and unless otherwise stated, conventional minimum mean square error (MMSE) combining with perfect channel state information (CSI) is assumed.
The simulation is conducted in an outdoor square area with 400 m sides. At the center, the base station with a 64-antenna uniform circular array is located at a 20 m height. The carrier frequency is 3.6 GHz with 20 MHz bandwidth. The noise temperature is 290 Kelvin and the NF is 5 dB. Path loss and LoS probability models stem from 3GPP UMi TS 36.814. The fading is Ricean with a 10 dB K-factor if LoS exists, otherwise it is Rayleigh. The 64 single-antenna repeaters/D-MIMO APs are distributed on a uniform grid at a 10 m height above the terminals. The terminals and repeaters have maximum 20 dBm output power. In addition, the repeaters are regulated by a maximum 70 dB amplification gain. The actual gain used at each repeater is determined as the maximum allowable amplification gain under these two constraints. To demonstrate the potential performance of RA-MIMO, it is assumed that the repeater-base station channels follow Rician fading with a LoS component. In practice, this can not always be guaranteed but is still likely with some deployment considerations.
Considering the additive noise at the base station and the repeaters, we point out that the combination of 1) received signal power variations across repeaters with 2) additive noise power at the repeaters may result in low-quality repeated signals that may deteriorate system performance (i.e., some repeaters act as noise sources). We address these challenges with a modified MMSE combiner to incorporate the amplified repeater noise stemming from repeater-base station channels into the matrix inversion of the linear MMSE filtering, in case of RA-MIMO.
Fig. 4 compares simulated cumulative distribution functions (CDFs) of the SINR for different systems, where the two MIMO architectures are compared to RA-MIMO systems with a 70 dB maximum amplification gain at each repeater. The performance of RA-MIMO systems with a 0 dB NF at each repeater is included as reference to depict an optimistic scenario with ideal-hardware repeaters. This enhancement in SINR signifies that RA-MIMO brings diversity and richer communication channels, thereby improving the low-performing terminals and bridging the performance gap between C-MIMO and D-MIMO systems. RA-MIMO systems can offer substantial benefits, but certain operational considerations must be taken into account. For instance, there is a risk for positive feedback loops between repeaters due to over-amplification if the amplificaiton control is not carefully designed, which can introduce instability in the system [6]. Additionally, the amplification process may also amplify noise, potentially affecting the overall signal quality. These factors should be carefully managed in practice to fully realize the advantages of RA-MIMO in improving network performance and efficiency.
@@
Figure 3: Simulated distributions of the SINR of all terminals for D-MIMO, C-MIMO, ideal-hardware RA-MIMO (0 dB NF) and realistic RA-MIMO (5 dB NF).
@@
Figure 4: Simulated distributions of the worst SINR among terminals per realization for D-MIMO, C-MIMO, ideal-hardware RA-MIMO (0 dB NF) and realistic RA-MIMO (5 dB NF).
@@
Shifting our focus to the worst-case SINR performance among terminals for each realization, this analysis provides deeper insights into the robustness and reliability of the MIMO configurations under the most challenging conditions. Fig. 4 presents the CDFs of the worst SINR among terminals for the different architectures. As depicted in the figure, RA-MIMO systems exhibit significant improvements over C-MIMO systems in terms of the worst communication performance, approaching the performance levels of D-MIMO. Together with the findings in Fig. 4, these results indicate that RA-MIMO systems have the potential to achieve a level of SINR performance uniformity that is comparable to that of D-MIMO.
@@ VI Acknowledgements
This work has been enabled by funding from the strategic research area ELLIIT, the KAW foundation, the Swedish Research Council, and H2020-REINDEER.

### v3

@@ Abstract
In what ways could cellular massive MIMO be improved? This technology has already been shown to bring huge performance gains. However, coverage holes and difficulties to transmit multiple streams to multi-antenna users because of insufficient channel rank remain issues. Distributed MIMO, also known as cell-free massive MIMO, might be the ultimate solution. However, while being a powerful technology, it is expensive to install backhaul, and it is a difficult problem to achieve accurate phase alignment for coherent multi-user beamforming on downlink. Another option is reflective intelligent surfaces – but they have large form factors and require a lot of training and control overhead, and probably, in practice, some form of active filtering to make them sufficiently band-selective.
We propose a new approach to densification of cellular systems, envisioning repeater-assisted cellular massive MIMO, where a large numbers of physically small and cheap wireless repeaters are deployed. They receive and retransmit signals instantaneously, appearing as active scatterers. Meaning that they appear as ordinary channel scatterers but with amplification. We elaborate on the requirements of such repeaters, show that the performance of these systems could potentially approach that of distributed MIMO, and outline future research directions.
@@ I Introduction
Massive multiple-input multiple-output (MIMO) has become the 5G technology and will continue to be one of the cornerstones in cellular systems. Sub-6 GHz time-division duplex (TDD) multi-user massive MIMO systems with reciprocity-based beamforming can serve many terminals in the same time-frequency resource, also in high mobility scenarios. By phase-coherent combining, an array gain is achieved, which can significantly improve signal quality and link budget.
@@
Although the opportunities with D-MIMO systems are many, they impose deployment issues. Many of the current solutions are not standardized and/or not yet sufficient in order to get the envisioned performance gains. Questions, such as how to get the data to the distributed units, where to perform the processing, how to synchronize the APs, and how to distribute the clocks and phase references still need further development. Furthermore, the main concepts in the standards are designed for co-located MIMO (C-MIMO) and going to D-MIMO would be a big step for industry to adopt. There would be a need for radical network architectural changes and a large increase in site and maintenance costs. In industry there can be some reluctance to big changes and smaller steps, each beneficial, would be more likely to be implemented.
To make the transition smoother and challenge the existing views on how densification of current cellular systems could be done in today’s sub-6 GHz systems, we envision repeater-assisted massive MIMO (RA-MIMO). A similar vision was presented in [1], but extra delay, increased delay spread and compromised channel reciprocity were identified as obstacles; these issues are solved in our vision. Repeaters have traditionally been deployed to improve coverage. Coverage improvements and capacity scaling in MIMO systems aided by repeaters was studied in LoS environments [2] and full-duplex radio relays were investigated to improve outdoor-to-indoor communication [3]. Today modern massive MIMO base stations enable multi-user MIMO operation. Challenging the current views on how densification could be done; we propose to rethink repeaters, could they be deployed inside the cell to support massive C-MIMO systems to increase macro diversity, acting as active scatterers, i.e. channel scatterers but with amplification, and potentially approach D-MIMO performance?
The use of the terms relay and repeater in literature is not always consistent. A repeater is here considered to be the same thing as a full-duplex relay, and instantaneously amplifies and re-transmits the signal, with a delay of at most a few hundred nanoseconds. However, a repeater is different from a half-duplex relay; the difference lies in the delay introduced. A half-duplex relay receives a signal of substantially longer duration (e.g., half a slot) and then either amplifies and forwards, or decodes and forwards, that signal. Repeaters (full-duplex relays) act as an active scatterer, but half-duplex relays do not.
@@
Other ways to enhance performance in sub-6 GHz systems, without creating extra cells, are to add extra APs or, use reflective intelligent surfaces (RISs). Table I presents a comparison between these three options; these technologies are of different maturity in terms of practical implementations. A RIS could perform the work of a repeater while bringing benefits stemming from the many unit cells. Network-controlled repeaters (NCRs) are a part of Release 18 [4, 5], while RISs has been decided to not become a study-item; several challenges need to be addressed before they can be used in practice [5].
Due to the different maturity of the technologies and the various possibilities to implement them, the power consumption numbers given are merely indicative. While sub-6 GHz low-power cellular repeater products on the market (not network-controlled) may consume as little as 10 W, an AP will consume significantly more due to the required processing capabilities and backhaul interface. The low power consumption potential of RIS yields a number at least an order of magnitude less than that of repeaters and APs. With radio frequency (RF) switch-based RIS elements, the power consumption will be dominated by the RIS control and switch driver circuitry [6], which may be further optimized and tailored for the application.
The advantages of deploying repeaters or RISs, compared to extra APs, are that neither phase synchronization nor backhaul is needed. While having these benefits in common, RISs have however drawbacks in comparison repeaters: it has a much larger form factor, requires significant control signaling, and incurs significant training and configuration overhead. Furthermore, a passive, or semi-passive RIS would not be nearly as band-selective as a repeater, which can be frequency-selective down to the level of a single channel. An active RIS, which could amplify signals, could act similar as a repeater but due to regulatory requirements it would likely require a large number of expensive components to filter the signal. As a first-order approximation, to achieve the same signal amplification as a repeater, a [passive] RIS would require a number of unit cells equal to the square root of the repeater (power) gain. To replace a 60-dB gain repeater, a RIS would need in the order of 31 × 31 ~ 1000 unit cells (for simplicity, we ignore the directivity of the patch antenna patterns here). Furthermore, to obtain the macro diversity, one would need many distributed units, which appears challenging with RISs. Challenges with incorporating RISs into cellular networks were discussed in [5], including a simulated comparison to NCRs, showing a performance gain in terms of spectral efficiency to the latter.
@@
Array gain Some High Some
@@ II Envisioning repeater-assisted massive MIMO
The signal received at the base station or the terminal will be a superposition of signals traveling via direct paths, including LoS and/or multipath reflections, and signals going via repeater(s). This means that the channels between a terminal and the base station is a result of several components: 1) the direct path between the terminal and base station, 2) paths between a terminal and a repeater, 3) paths between a repeater and a scatterer, 4) paths between repeaters, 5) paths between a repeater and the base station, and 6) paths between a scatterer and the base station. As this is only concatenations of linear time-invariant (LTI) systems, the result is also an LTI system. Its response is (in principle) the same in uplink and downlink, enabling reciprocity-based multi-user beamforming. Repeaters should act as active scatterers in the environment, i.e. similar to ordinary channel scatterers, but with higher gain. The repeaters will naturally introduce a delay, i.e. a frequency-dependent phase shift, which is typically not a problem as it is comparable to the frequency response of a far-away scatterer.
III Repeaters as active channel scatterers
@@ III-A Requirements on the repeater
To support the vision of RA-MIMO systems, the key aspect is that it must be possible to realize repeaters that have the same properties as channel scatterers present in the propagation environment. With the same properties, they would be a transparent part from both the base station and terminal perspective. To act as a channel scatterer, the repeater must satisfy the following requirements:
@@
    Noise is unavoidable and stems from multiple sources: foremost thermal noise and phase noise. There is a limit on how noisy a repeater can be in order to not act as a limiting noise source.
@@ III-C Dual-antenna repeaters
RF band filters can be used, see 3) in Fig. 2, to attenuate out-of-band signals before the LNA, and to comply with out-of-band emission requirements. However, such filters may not be sufficient to fulfill gain requirements outside the repeater passband. Better selectivity can be achieved by baseband filtering, as shown in 4) in Fig. 2: the received signal is then down-converted to baseband, filtered, and up-converted using the same local oscillator (LO).
@@ IV Performance of repeater-assisted cellular massive MIMO
After elaborating on how to implement repeaters to act as active scatterers we here aim to show the potential of RA-MIMO, comparing it to massive C-MIMO, small cells, and densely deployed D-MIMO, which serves as a benchmark for ideal performance, e.g. assuming no synchronization issues or other impairments. We consider multi-user MIMO transmission, where either a massive MIMO array located at the base station (first two cases) or distributed APs (last two cases) receive uplink data from multiple terminals, whereas the repeaters in RA-MIMO are located at the same positions as the APs in the last two cases. For simplicity, the results disregard interactions between repeaters, however, the maximum amplification gain is chosen such that the system becomes stable without positive feedback (using the criteria in [8]). The signal-to-interference-plus-noise ratio (SINR) after the uplink combiner is used as performance metric, with conventional minimum mean square error (MMSE) combining and perfect channel state information (CSI) assumed unless otherwise stated. Cumulative distribution functions (CDFs) based on both channel distributions and user locations are generated, using 100,000 realizations. The terminals are randomly positioned in each realization. For the AP selection in the small cell scenario, a greedy strategy leveraging the Hungarian algorithm to optimize total signal strength across all terminals is used.
@@
Figure 3: Simulated SINR distributions for all terminals, for D-MIMO, C-MIMO, ideal-hardware RA-MIMO (0 dB NF), realistic RA-MIMO (5 dB NF), and small cells.
@@
Figure 4: Simulated distributions of the worst SINR among all terminals per realization, for the same systems as in Fig. 4.
The simulation is conducted in an outdoor square area with 400 m sides. At the center, the base station with a 64-antenna uniform circular array is located at a 20 m height. The carrier frequency is 3.6 GHz with 20 MHz bandwidth. The noise temperature is 290 Kelvin and the NF is 5 dB. Path loss and LoS probability models stem from 3GPP UMi TS 36.814. The fading is Ricean with a 10 dB K-factor if LoS exists, otherwise it is Rayleigh. The 64 single-antenna repeaters/APs are distributed on a uniform grid at a 10 m height above the terminals. The terminals and repeaters have maximum 20 dBm output power. In addition, the repeaters are regulated by a maximum 70 dB amplification gain. The actual gain used at each repeater is determined as the maximum allowable amplification gain under these two constraints. To demonstrate the potential performance of RA-MIMO, it is assumed that the repeater-base station channels follow Rician fading with a LoS component. In practice, this can not always be guaranteed but is still likely with some deployment considerations.
Considering the additive noise at the base station and the repeaters, we note that the combination of 1) received signal power variations across repeaters with 2) additive noise power at the repeaters may result in low-quality repeated signals that may deteriorate system performance (i.e., some repeaters act as noise sources). We address these challenges with a modified MMSE combiner to incorporate the amplified repeater noise stemming from repeater-base station channels into the matrix inversion of the linear MMSE filtering, in case of RA-MIMO.
Fig. 4 compares simulated CDFs of the SINR for different systems, where the two MIMO architectures are compared to RA-MIMO systems with a 70 dB maximum amplification gain at each repeater. The performance of RA-MIMO systems with a 0 dB NF at each repeater is included as reference to depict an optimistic scenario with ideal-hardware repeaters. This enhancement in SINR signifies that RA-MIMO brings diversity and richer communication channels, thereby improving the low-performing terminals and bridging the performance gap between C-MIMO and D-MIMO systems. RA-MIMO systems can offer substantial benefits, but certain operational considerations must be taken into account. For instance, there is a risk for positive feedback loops between repeaters due to over-amplification if the amplificaiton control is not carefully designed, which can introduce instability in the system [8]. Additionally, the amplification process may also amplify noise, affecting the signal quality. These factors should be carefully managed in practice to fully realize the advantages of RA-MIMO.
@@
Shifting our focus to the worst-case SINR performance among terminals for each realization, this analysis provides insights into the robustness and reliability of the MIMO configurations under challenging conditions. Fig. 4 presents the CDFs of the worst SINR among terminals for the different architectures. As depicted, RA-MIMO systems exhibit significant improvements over C-MIMO systems in terms of the worst communication performance, approaching the performance levels of D-MIMO. Together with the findings in Fig. 4, these results indicate that RA-MIMO systems have the potential to achieve a level of SINR performance uniformity that is comparable to that of D-MIMO.
@@ VI Acknowledgements
This work was supported by the strategic research area ELLIIT, the KAW foundation, the Swedish Research Council, and H2020-REINDEER.